OpenStudy (anonymous):

Simplify: sqrt(c) (3 + 2sqrt(5))

OpenStudy (anonymous):

\[\sqrt{c} (3+2\sqrt{5})\] That's a better way of putting it.

OpenStudy (tkhunny):

Who says it can be simplified?

OpenStudy (anonymous):

My test

OpenStudy (tkhunny):

Well, I would have to disagree with your test, but \(3\sqrt{6} + 2\sqrt{5c}\) might be considered "simplified" - based on some weird, twisted definition of "simplified".

OpenStudy (anonymous):

Well my test, it is on Virtual School. and i have to distribute apparently

OpenStudy (tkhunny):

Fair enough. Just keep in mind that there is NO clear or widely-accepted definition of "simplified".